OpenStudy (anonymous):

Find the slope-predictor formula and write an equation for the line tangent to (a, f(a)) for f(x) = -2x^2+3x-7

OpenStudy (anonymous):

is that what they are calling derivatives now? slope-predictor formula??

OpenStudy (anonymous):

\[f(x)=-2x^2+3x-7\] \[f'(x)=-4x+3\] so \[f'(a)=-4a+3\] and that his the slope

OpenStudy (anonymous):

then use point - slope formula to find the equation of the line: \[y-y_1=m(x-x_1)\] with \[x_1=a, y_1=f(a), m=(-4a+3)\]

OpenStudy (anonymous):

you get \[y-f(a)=(-4a+3)(x-a)\] and then some algebra

OpenStudy (anonymous):

are you perhaps supposed to find \[f'(a)\] by hand using \[\lim_{x\rightarrow a}\frac{f(a)-f(x)}{a-x}\]?

OpenStudy (anonymous):

yes